# Regge trajectories for the triply heavy bottom-charm baryons in the diquark picture

### v1

@@ I Introduction
In the diquark picture, the triply heavy bottom-charm baryons, $\Omega_{ccb}$ and $\Omega_{cbb}$, are composed of one doubly diquark ($(cc)$ or $(bb)$) and one heavy quark ($b$ or $c$). The triply heavy bottom-charm baryons have been studied by a vast variety of approaches including the lattice QCD Mathur:2018epb , the renormalization group procedure for effective particles Serafin:2018aih , the relativistic quark model based on the quark-diquark picture in the quasipotential approach Faustov:2021qqf , the non-relativistic quark model Silvestre-Brac:1996myf ; Roberts:2007ni ; deArenaza:2024dhe , the constituent quark model Yang:2019lsg , the bag model Hasenfratz:1980ka , the sum rules Zhang:2009re ; Najjar:2024deh ; Wang:2011ae , the relativistic three-quark model Martynenko:2007je , the variational method Jia:2006gw , and baryon Faddeev equations Yin:2019bxe .
The triply heavy bottom-charm baryons are also discussed by the Regge trajectories  [^1] A Regge trajectory of hadrons is assumed to be written as $M=m_{R}+\beta_{x}(x+c_{0})^{\nu}$ $(x=l,\,n_{r})$, where $M$ is mass of the bound state, $l$ is the the orbital angular momentum, and $n_{r}$ is the radial quantum number. $m_{R}$ and $\beta_{x}$ are parameters. For simplicity, the plots in the $(M,\,x)$ plane Chen:2022flh ; Chen:2023cws , in the $(M,\,(x-c_{0})^{\nu})$ plane Burns:2010qq , in the $(M^{2},\,x)$ plane Chen:2018nnr ; Chen:2021kfw , in the $((M-m_{R})^{2},\,x)$ plane Chen:2023djq ; Chen:2023web or in the $((M-m_{R})^{1/{\nu}},\,x)$ plane Xie:2024dfe are all called the Chew-Frautschi plots. The Regge trajectories can be plotted in these different planes. Ishida:1994pf ; Chen:2023djq ; Oudichhya:2023pkg . In preceding works, the Regge trajectories for the triply heavy bottom-charm baryons are mostly the $\lambda$-trajectories, with few studies addressing the $\rho$-trajectories Ishida:1994pf . In this work, we investigate both the $\lambda$-trajectories and the $\rho$-trajectories for the triply heavy bottom-charm baryons. The Regge trajectories take different forms in different energy regions Chen:2022flh ; Chen:2021kfw . In the case of the $\lambda$-mode, the triply heavy bottom-charm baryons are the heavy-heavy systems because both the diquark ($(bb)$ or $(cc)$) and the quark ($b$ or $c$) are heavy. In the case of the $\rho$-mode, the triply heavy bottom-charm baryons also show the heavy-heavy properties because the diquarks $(bb)$ and $(cc)$ are the heavy-heavy systems. By employing the diquark Regge trajectory relation Feng:2023txx , we present the Regge trajectory relations (II.5) for the triply heavy bottom-charm baryons. The masses of the $\lambda$-excited states and the $\rho$-excited states are estimated, and both the $\lambda$-trajectories and the $\rho$-trajectories are discussed.
@@
II Regge trajectory behaviors for baryons
In this section, we discuss the behaviors of the Regge trajectories for various baryons with the help of the diquark Regge trajectories Chen:2023cws ; Feng:2023txx ; Chen:2023ngj . The behaviors of the Regge trajectories for various tetraquarks are referred to Ref. plan .
@@ II.2 Spinless Salpeter equation
II.3 Regge trajectory relations
@@
By employing the Bohr-Sommerfeld quantization approach Brau:2000st , the parameterized formula can be written as Chen:2022flh ; Chen:2021kfw
@@
Eq. (15) with (17) is obtained in the limit $m_{1}\to\infty$ and $m_{2}\to 0$. There are different ways to include the light constituent’s mass Selem:2006nd ; Nielsen:2018uyn ; Sonnenschein:2018fph ; MartinContreras:2020cyg ; Chen:2023cws ; Chen:2023ngj ; Chen:2022flh ; Chen:2014nyo ; Afonin:2014nya ; Sergeenko:1994ck . Two modified formulas are proposed in Ref. Chen:2023cws , which can universally describe both the heavy-light mesons and the heavy-light diquarks. One is Eq. (15) with (9). Another reads
@@
where $\beta_{x}$ is in (24). Eq. (15) with (9) is an extension of Afonin:2014nya
@@
while (18) with (19) is based on the results in Selem:2006nd ; Sonnenschein:2018fph . Although they give different behavior of $m_{2}$, Eq. (15) with (9) and Eq. (18) with (19) produce consistent results for $l,\,n_{r}<10$ and have the same behavior $M{\sim}x^{1/2}$ Chen:2023cws .
@@
By employing the Bohr-Sommerfeld quantization approach Brau:2000st , the parameterized formula can be written as Chen:2022flh ; Chen:2021kfw
@@
When Eq. (15) with (9) is employed to discuss the heavy-light systems, and Eq. (26) with (9) is used to discuss the light systems, summarising Eqs. (8), (15) with (9) and (26) with (9), we have a general form of the Regge trajectories Chen:2022flh ; Xie:2024dfe
@@
II.4 Behaviors of the Regge trajectories for baryons
As an example, we consider the baryon $(qq^{\prime}){{q^{\prime\prime}}}$, in which the light diquark $(qq^{\prime})$ is composed of two light quarks and ${{q^{\prime\prime}}}$ is also light. When discussing the Regge trajectories of the light diquarks, Eq. (25) gives the same behavior as Eq. (26) with (9),
@@ II.4 Behaviors of the Regge trajectories for baryons
where $M_{\rho}$ is the diquark mass. Generally speaking, the diquark Regge trajectories are not the same as the $\rho$-trajectories of baryons. For a light diquark composed of two light quarks, the behavior of the diquark Regge trajectory is definite (see Eq. (29)) if the masses of light quarks are not considered. However, it is indefinite when discussing the $\rho$- and $\lambda$-trajectories of baryons composed of the light diquark and/or the light quark. If the diquark $(qq^{\prime})$ is regarded as being light, when Eq. (25) is chosen as the $\lambda$-trajectory relation, it will give $M{\sim}M_{\rho}^{3/2}$ while Eq. (26) with (9) gives $M{\sim}M_{\rho}$ due to different ways to include the light constituent’s mass $M_{\rho}$ [corresponding to $m_{1,2}$ in Eq. (25) and (26) with (9).]. They give two different behaviors of the $\rho$-trajectories. Using Eq. (29), Eq. (25) gives $M{\sim}x_{\rho}^{3/4}$ while Eq. (26) with (9) gives $M{\sim}x_{\rho}^{1/2}$. When discussing the $\rho$-trajectories, highly excited diquark will be involved. The first radially and orbitally excited $(ud)$ are about $1.3$ GeV. The masses of the $1^{1}p_{1}$ states of $(us)$ and $(ss)$ are much heavier, $1.46$ GeV and $1.65$ GeV Chen:2023ngj , respectively. They approximate or even exceed the mass of the charm quark. Therefore, the light diquark composed of two light quarks can be regarded as being heavy when discussing the $\rho$-trajectories of baryons. Then, Eq. (15) with (9) or Eq. (18) with (19) can be chosen as the ${\lambda}$-trajectory formula. They will give the same $\rho$-trajectory behavior, $M{\sim}M_{\rho}$. Using Eq. (29), we get $M{\sim}x_{\rho}^{1/2}$.
@@ 5 $(QQ^{\prime})q$ $x_{\rho}^{2/3}$ $x_{\lambda}^{1/2}$ $x_{\rho}^{2/3}$ $x_{\lambda}^{1/2}$
If the light diquark is regarded as being light, $\lambda$-trajectories behave as $M{\sim}x^{1/2}_{\lambda}$ no matter if Eq. (25) or Eq. (26) with (9) is chosen as the $\lambda$-trajectory formula. If the light diquark is regarded as being heavy, $\lambda$-trajectories also behave as $M{\sim}x_{\lambda}^{1/2}$ according to Eq. (15) with (9) or Eq. (18) with (19). Therefore, the behavior of the $\lambda$-trajectories is definite whether the light diquark is regarded as being light or being heavy.
@@
When the diquark is doubly heavy or heavy-light, it is clear how to introduce its large mass, and the way of introducing the light mass does not affect the Regge trajectory behaviors (see Eqs. (8) with (9) and Eq. (15) with (9) or Eq. (18) with (19)). The Regge trajectory formulas can be written explicitly, and then the behaviors of the $\lambda$-trajectory and the $\rho$-trajectory are definite. For example, the Regge trajectory for the baryons composed of one doubly heavy diquark and one heavy quark can be written explicitly as Eq. (II.5) with (II.5). Other types of baryons can be discussed in a similar manner. The Regge trajectory behaviors of different types of baryons are listed in Table 3.
@@ II.5 Regge trajectory relations for the triply heavy baryons
In Eq. (II.5), $M$ is the baryon mass, and $M_{\rho}$ is the diquark mass. The second relation in Eq. (II.5) is used to calculate the diquark masses. The first relation in Eq. (II.5) together with other relations in Eq. (II.5) and (II.5) is employed to calculate the baryon mass. It is obvious that there are two series of Regge trajectories for the triply heavy baryons.
@@ III.1 Parameters
The quark masses, the string tension $\sigma$, and the parameter $C$ are from Ref. Faustov:2021qqf . The parameters for the doubly heavy diquarks $(cc)$ and $(bb)$ are from Ref. Feng:2023txx and listed in Table 4. With these parameters determined, the $\rho-$modes and the diquark masses can be discussed (see Eqs. (II.5) and (II.5)). To discuss the $\lambda-$modes excited states, the parameters $c_{fL}$, $c_{fN_{r}}$, $c_{0{L}}$, and $c_{0N_{r}}$ should be determined (see Eqs. (32) and (33)). More detailed discussions are in Xie:2024dfe .
@@ III.2 $\lambda-$ and $\rho-$trajectories
Table 5: The spin-averaged masses of the $\lambda$-excited states of $\Omega_{ccb}$ and $\Omega_{cbb}$ (in ${{\rm GeV}}$). The notation in Eq. (1) is rewritten as $|n^{2s_{d}+1}l_{j_{d}},N^{2j+1}L_{J}\rangle$. And $|n^{2s_{d}+1}l_{j_{d}},NL\rangle$ denotes the spin-averaged states. Eqs. (II.5), (32) and (33) are used.
@@
Baryon $|n^{2s_{d}+1}l_{j_{d}},NL\rangle$ Our Ref. Faustov:2021qqf Ref. Silvestre-Brac:1996myf Ref. Yang:2019lsg Ref. Roberts:2007ni
$\Omega_{ccb}$ $|1^{3}s_{1},1S\rangle$ 7.88 7.99 8.04 8.02 8.26
@@
Figure 2: The $\lambda$-trajectories for the triply heavy bottom-charm baryons. Circles represent the predicted data and the black lines are the $\lambda$-trajectories, see Eq. (II.5). Data are listed in Table 5.
Figure 3: The $\rho$-trajectories for the triply heavy bottom-charm baryons. Circles represent the predicted data and the black lines are the $\rho$-trajectories, see Eq. (II.5). Data are listed in Table 6.
By using Eqs. (II.5), (32) and (33), along with the parameters in Table 4, the spin-averaged masses of the $\lambda$-excitations and $\rho$-excitations are estimated (see Tables 5 and 6). The obtained results are in agreement with other theoretical predictions, see Table 7. The $\lambda$-trajectories and $\rho$-trajectories are shown in Figs. 2 and 3.
@@
Both the diquarks ($(bb)$ and $(cc)$) and the quarks ($b$ and $c$) are heavy; therefore, the case of the triply heavy bottom-charm baryons is simple, and the Regge trajectory relations can be written explicitly, see (II.5). From Eq. (II.5), we can see easily that the $\lambda$-trajectories and $\rho$-trajectories have the same behaviors, $M{\sim}x^{2/3}_{\lambda}$ and $M{\sim}x^{2/3}_{\rho}$.
The curvature of the Regge trajectories holds significant importance Chen:2018bbr . In potential models, the curvature is related to the dynamic equation and the confining potential. In this work, we show that for the triply heavy bottom-charm baryon, not only the $\lambda$-trajectories but also the $\rho$-trajectories are concave in the $(M^{2},\,x)$ planes, see Eq. (II.5).

### v3

@@ I Introduction
In the diquark picture, the triply heavy bottom-charm baryons, $\Omega_{ccb}$ and $\Omega_{cbb}$, are composed of one doubly diquark ($(cc)$ or $(bb)$) and one heavy quark ($b$ or $c$). The triply heavy bottom-charm baryons have been studied by a vast variety of approaches including the lattice QCD Mathur:2018epb ; Brown:2014ena , the renormalization group procedure for effective particles Serafin:2018aih , the relativistic quark model based on the quark-diquark picture in the quasipotential approach Faustov:2021qqf , the non-relativistic quark model Silvestre-Brac:1996myf ; Roberts:2007ni ; deArenaza:2024dhe , the constituent quark model Yang:2019lsg , the bag model Hasenfratz:1980ka , the sum rules Zhang:2009re ; Najjar:2024deh ; Wang:2011ae , the relativistic three-quark model Martynenko:2007je , the variational method Jia:2006gw , and baryon Faddeev equations Yin:2019bxe .
The triply heavy bottom-charm baryons are also discussed by the Regge trajectories  [^1] A Regge trajectory of hadrons is assumed to be written as $M=m_{R}+\beta_{x}(x+c_{0})^{\nu}$ $(x=l,\,n_{r})$, where $M$ is mass of the bound state, $l$ is the the orbital angular momentum, and $n_{r}$ is the radial quantum number. $m_{R}$ and $\beta_{x}$ are parameters. For simplicity, the plots in the $(M,\,x)$ plane Chen:2022flh ; Chen:2023cws , in the $(M,\,(x-c_{0})^{\nu})$ plane Burns:2010qq , in the $(M^{2},\,x)$ plane Chen:2018nnr ; Chen:2021kfw , in the $((M-m_{R})^{2},\,x)$ plane Chen:2023djq ; Chen:2023web or in the $((M-m_{R})^{1/{\nu}},\,x)$ plane Xie:2024dfe are all called the Chew-Frautschi plots. The Regge trajectories can be plotted in these different planes. Ishida:1994pf ; Chen:2023djq ; Oudichhya:2023pkg . In preceding works, the Regge trajectories for the triply heavy bottom-charm baryons are mostly the $\lambda$-trajectories, with few studies addressing the $\rho$-trajectories Ishida:1994pf . In this work, we investigate both the $\lambda$-trajectories and the $\rho$-trajectories for the triply heavy bottom-charm baryons. The Regge trajectories take different forms in different energy regions Chen:2022flh ; Chen:2021kfw . In the case of the $\lambda$-mode, the triply heavy bottom-charm baryons are the heavy-heavy systems because both the diquark ($(bb)$ or $(cc)$) and the quark ($b$ or $c$) are heavy. In the case of the $\rho$-mode, the triply heavy bottom-charm baryons also show the heavy-heavy properties because the diquarks $(bb)$ and $(cc)$ are the heavy-heavy systems. By employing the diquark Regge trajectory relation Feng:2023txx , we present the Regge trajectory relations (II.5) and (II.5) for the triply heavy bottom-charm baryons. The masses of the $\lambda$-excited states and the $\rho$-excited states are estimated, and both the $\lambda$-trajectories and the $\rho$-trajectories are discussed.
@@
II Regge trajectory behaviors for various baryons
The Regge trajectories for different systems take different forms Chen:2022flh ; Chen:2021kfw . In this section, we use the diquark Regge trajectories Chen:2023cws ; Feng:2023txx ; Chen:2023ngj to discuss the behaviors of the baryon Regge trajectories. We show that the behaviors of the Regge trajectories for the triply heavy baryons are different from the behaviors of the Regge trajectories for other types of baryons. The behaviors of the Regge trajectories for various tetraquarks are referred to Ref. Xie:2024dfe .
@@ II.2 Spinless Salpeter equation
II.3 Regge trajectory relations for different systems
In this subsection, we present the Regge trajectory relations for the heavy-heavy systems, the heavy-light systems, and the light systems.
@@ II.3 Regge trajectory relations for different systems
By employing the Bohr-Sommerfeld quantization approach Brau:2000st , the parameterized formula can be obtained and is written as Chen:2022flh ; Chen:2021kfw
@@
Eqs. (15) and (17) are obtained in the limit $m_{1}\to\infty$ and $m_{2}\to 0$. There are different ways to include the light constituent’s mass Selem:2006nd ; Nielsen:2018uyn ; Sonnenschein:2018fph ; MartinContreras:2020cyg ; Chen:2023cws ; Chen:2023ngj ; Chen:2022flh ; Chen:2014nyo ; Afonin:2014nya ; Sergeenko:1994ck . Two modified formulas are proposed in Ref. Chen:2023cws , which can universally describe both the heavy-light mesons and the heavy-light diquarks. One is Eq. (15) together with Eq. (9). Another reads
@@
where $\beta_{x}$ is in (24). Eqs. (15) and (9) are the extension of Afonin:2014nya
@@
while (18) and (19) are based on the results in Selem:2006nd ; Sonnenschein:2018fph . Although they give different behavior of $m_{2}$, Eqs. (15), (9) , (18) and (19) produce consistent results for $l,\,n_{r}<10$ and have the same behavior $M{\sim}x^{1/2}$ Chen:2023cws .
@@
By employing the Bohr-Sommerfeld quantization approach Brau:2000st , the parameterized formula can be obtained, which is written as Chen:2022flh ; Chen:2021kfw
@@
When Eqs. (8) and (9) are applied to discuss the heavy-heavy systems, Eqs. (15) and (9) are employed to discuss the heavy-light systems, and Eqs. (26) and (9) are used to discuss the light systems, summarising Eqs. (8), (9), (15) and (26), we have a general form of the Regge trajectories Chen:2022flh ; Xie:2024dfe
@@
II.4 Behaviors of the Regge trajectories for various baryons
In this subsection, we apply the results obtained in the preceding subsection to discuss the behaviors of the Regge trajectories for various baryons. We find that different from the behaviors of the Regge trajectories for other types of baryons, the $\lambda$-trajectories and the $\rho$-trajectories for the triply heavy baryons have the same behaviors, $M{\sim}x^{2/3}_{\lambda}$ and $M{\sim}x^{2/3}_{\rho}$ or $M^{2}{\sim}x^{2/3}_{\lambda}$ and $M^{2}{\sim}x^{2/3}_{\rho}$, see Table 3. Moreover, the Regge trajectory relations for the triply heavy baryons can be written explicitly, see Eqs. (II.5), (II.5) and (II.5).
As an example, we consider the baryon $(qq^{\prime}){{q^{\prime\prime}}}$, in which the light diquark $(qq^{\prime})$ is composed of two light quarks and ${{q^{\prime\prime}}}$ is also light. When discussing the Regge trajectories of the light diquarks, Eq. (25) gives the same behavior as Eqs. (26) and (9),
@@ II.4 Behaviors of the Regge trajectories for various baryons
where $M_{\rho}$ is the diquark mass. Generally speaking, the diquark Regge trajectories are not the same as the $\rho$-trajectories of baryons. For a light diquark composed of two light quarks, the behavior of the diquark Regge trajectory is definite (see Eq. (29)) if the masses of light quarks are not considered. However, it is indefinite when discussing the $\rho$- and $\lambda$-trajectories of baryons composed of the light diquark and/or the light quark. If the diquark $(qq^{\prime})$ is regarded as being light, when Eq. (25) is chosen as the $\lambda$-trajectory relation, it will give $M{\sim}M_{\rho}^{3/2}$ while Eqs. (26) and (9) gives $M{\sim}M_{\rho}$ due to different ways to include the light constituent’s mass $M_{\rho}$ [corresponding to $m_{1,2}$ in Eqs. (25), (26) and (9).]. They give two different behaviors of the $\rho$-trajectories. Using Eq. (29), Eq. (25) gives $M{\sim}x_{\rho}^{3/4}$ while Eqs. (26) and (9) give $M{\sim}x_{\rho}^{1/2}$. When discussing the $\rho$-trajectories, highly excited diquark will be involved. The first radially and orbitally excited $(ud)$ are about $1.3$ GeV. The masses of the $1^{1}p_{1}$ states of $(us)$ and $(ss)$ are much heavier, $1.46$ GeV and $1.65$ GeV Chen:2023ngj , respectively. They approximate or even exceed the mass of the charm quark. Therefore, the light diquark composed of two light quarks can be regarded as being heavy when discussing the $\rho$-trajectories of baryons. Then, Eqs. (15) and (9) or Eqs. (18) and (19) can be chosen as the ${\lambda}$-trajectory formula. They will give the same $\rho$-trajectory behavior, $M{\sim}M_{\rho}$. Using Eq. (29), we get $M{\sim}x_{\rho}^{1/2}$.
@@ 5 $(QQ^{\prime})q$ $x_{\rho}^{2/3}$ $x_{\lambda}^{1/2}$ $x_{\rho}^{2/3}$ $x_{\lambda}^{1/2}$
If the light diquark is regarded as being light, $\lambda$-trajectories behave as $M{\sim}x^{1/2}_{\lambda}$ no matter if Eq. (25) or Eqs. (26) and (9) are chosen as the $\lambda$-trajectory formula. If the light diquark is regarded as being heavy, $\lambda$-trajectories also behave as $M{\sim}x_{\lambda}^{1/2}$ according to Eqs. (15) and (9) or Eqs. (18) and (19). Therefore, the behavior of the $\lambda$-trajectories is definite whether the light diquark is regarded as being light or being heavy.
@@
When the diquark is doubly heavy or heavy-light, it is clear how to introduce its large mass, and the way of introducing the light mass does not affect the Regge trajectory behaviors (see Eqs. (8), (9), (15), (18) and (19)). The Regge trajectory formulas can be written explicitly, and then the behaviors of the $\lambda$-trajectory and the $\rho$-trajectory are definite. For example, the Regge trajectory for the baryons composed of one doubly heavy diquark and one heavy quark can be written explicitly as Eqs. (II.5), (II.5) and (II.5). Other types of baryons can be discussed in a similar manner. The Regge trajectory behaviors of different types of baryons are listed in Table 3.
@@ II.5 Regge trajectory relations for the triply heavy baryons
In Eq. (II.5), $M$ is the mass of the triply heavy baryon, and $M_{\rho}$ is the mass of the doubly heavy diquark. The second relation in Eq. (II.5) is used to calculate the diquark masses. The first relation in Eq. (II.5) and other relations in Eqs. (II.5) and (II.5) are employed to calculate the masses of the triply heavy bottom-charm baryons.
According to Eqs. (II.5) and (II.5), we have
$M=M_{\rho}+m_{Q^{\prime\prime}}+C+\beta_{x_{\lambda}}(x_{\lambda}+c_{0x_{\lambda}})^{2/3}$ (32)
when the diquark is regarded as a constituent and the structure of it is not considered. When the diquark is considered as a bound state composed of two heavy quarks Feng:2023txx , we have
$M=$ $m_{Q}+m_{Q^{\prime}}+m_{Q^{\prime\prime}}+\frac{3}{2}C$
$+\beta_{x_{\lambda}}(x_{\lambda}+c_{0x_{\lambda}})^{2/3}+\beta_{x_{\rho}}(x_{\rho}+c_{0x_{\rho}})^{2/3}$ (33)
from Eqs. (II.5) and (II.5). We can see from Eq. (II.5) that there are two series of Regge trajectories for the triply heavy baryons: the $\lambda$-trajectories and the $\rho$-trajectories. The Regge trajectory relations (Eqs. (II.5), (II.5) and (II.5)) for the triply heavy baryons have the similar forms as the Regge trajectory relations for the triply heavy triquarks Song:2024bkj .
@@ III.1 Parameters
The quark masses, the string tension $\sigma$, and the parameter $C$ are from Ref. Faustov:2021qqf . The parameters for the doubly heavy diquarks $(cc)$ and $(bb)$ are from Ref. Feng:2023txx and listed in Table 4. With these parameters determined, the $\rho-$modes and the diquark masses can be discussed (see Eqs. (II.5), (II.5) and (II.5)). To discuss the $\lambda-$modes excited states, the parameters $c_{fL}$, $c_{fN_{r}}$, $c_{0{L}}$, and $c_{0N_{r}}$ should be determined (see Eqs. (34) and (35)). More detailed discussions are in Xie:2024dfe .
@@ III.2 $\lambda-$ and $\rho-$trajectories
Table 5: The spin-averaged masses of the $\lambda$-excited states of $\Omega_{ccb}$ and $\Omega_{cbb}$ (in ${{\rm GeV}}$). The notation in Eq. (1) is rewritten as $|n^{2s_{d}+1}l_{j_{d}},N^{2j+1}L_{J}\rangle$. And $|n^{2s_{d}+1}l_{j_{d}},NL\rangle$ denotes the spin-averaged states. Eqs. (II.5) (or (II.5)), (34) and (35) are used.
@@
Baryon $|n^{2s_{d}+1}l_{j_{d}},NL\rangle$ Our Ref. Faustov:2021qqf Ref. Silvestre-Brac:1996myf Ref. Yang:2019lsg Ref. Roberts:2007ni Mathur:2018epb Brown:2014ena
$\Omega_{ccb}$ $|1^{3}s_{1},1S\rangle$ 7.88 7.99 8.04 8.02 8.26 8.02 8.03
@@
Figure 2: The $\lambda$-trajectories for the triply heavy bottom-charm baryons. Circles represent the predicted data and the black lines are the $\lambda$-trajectories, see Eq. (II.5) or (II.5). Data are listed in Table 5.
Figure 3: The $\rho$-trajectories for the triply heavy bottom-charm baryons. Circles represent the predicted data and the black lines are the $\rho$-trajectories, see Eq. (II.5) or (II.5). Data are listed in Table 6.
By using Eqs. (II.5) (or (II.5)), (II.5), (34) and (35), along with the parameters in Table 4, the spin-averaged masses of the $\lambda$-excitations and $\rho$-excitations are estimated (see Tables 5 and 6). The obtained results are in agreement with other theoretical predictions, see Table 7. The $\lambda$-trajectories and $\rho$-trajectories are shown in Figs. 2 and 3.
@@
Both the diquarks ($(bb)$ and $(cc)$) and the quarks ($b$ and $c$) are heavy; therefore, the case of the triply heavy bottom-charm baryons is simple, and the Regge trajectory relations can be written explicitly, see (II.5) and (II.5). From Eq. (II.5) or (II.5), we can see easily that the $\lambda$-trajectories and $\rho$-trajectories have the same behaviors, $M{\sim}x^{2/3}_{\lambda}$ and $M{\sim}x^{2/3}_{\rho}$.
The curvature of the Regge trajectories holds significant importance Chen:2018bbr . In potential models, the curvature is related to the dynamic equation and the confining potential. In this work, we show that for the triply heavy bottom-charm baryon, not only the $\lambda$-trajectories but also the $\rho$-trajectories are concave in the $(M^{2},\,x)$ planes, see Eq. (II.5) or (II.5).